OpenStudy (anonymous):

1.) write the sum using summation notation assuming the suggested pattern continues 4 - 12 +36 -108 +...

OpenStudy (anonymous):

@welshfella @amoodarya

OpenStudy (anonymous):

@Michele_Laino @iambatman @ganeshie8

OpenStudy (anonymous):

@Astrophysics

OpenStudy (michele_laino):

Hint: it is a geometric sequence, whose constant is: q=-108/36 = 36/(-12) = (-12/4= -3

OpenStudy (michele_laino):

and first term is a1=4

OpenStudy (anonymous):

4(-3)^n or 4(-3)^(n-1)

OpenStudy (michele_laino):

the n-th term is therefore: \[\Large {a_n} = {a_1}{q^{n - 1}} = 4 \times {\left( { - 3} \right)^{n - 1}}\]

OpenStudy (michele_laino):

since we start to count, from n=1

OpenStudy (anonymous):

thanks @Michele_Laino

OpenStudy (michele_laino):

:)